# Probing exotic long-lived particles from the prompt side using the CONTUR method

### v2

@@ 3 Adapting Contur to LLP topologies
We propose the following approach to treat LLPs in the Rivet and Contur machinery. The basic idea is to evaluate, for a given set of model parameters, the lab-frame decay lengths for all particles with small widths (and therefore long lifetimes) as determined by the event generator. We can then sort the particles into three categories: a) those which decayed early enough to be treated as prompt, b) those which decayed within the detector volume and hence would be unlikely to be picked up by the standard reconstruction algorithms used for measurements if not actively discarded, and c) those which decayed outside the bounds of the detector and which would, if neutral, be reconstructed as missing transverse energy. The individual particles which were not reconstructible can be manually removed from the simulated events (for instance with PyHEPMC BUCKLEY2021107310 ), while events which contained decays within the detector volume, which may cause signatures similar to instrumental noise, can be entirely removed. Once the relevant particles and/or events have been removed from the simulations, the rest of the Contur workflow can then run normally. We have implemented this workflow, and extracted new constraints on LLP models as discussed in Sections 456 and 7. This approach is conceptually related to reweighting-based approaches previously described in SModelS Heisig:2018kfq and in a standalone study of dark radiation using LLP searches Bernreuther:2022bdw , however, it goes further by making event-by-event decisions on which particles to reconstruct, and it is the first time that any such approach has been used with detector-corrected LHC measurements. We will now review the step-by-step implementation of our method, including technical details.
@@
We determine a transverse “prompt threshold” $p_{th}$ before which particles can be considered indistinguishable from prompt. Indeed, if the decay of the LLP takes place before $p_{th}$, then tracks and deposits from the decay products should be reconstructible using the standard ATLAS and CMS algorithms. To choose the value of the threshold, we are guided by Refs. ATLAS:2023nze and CMS:2014pgm . The ultimate limiting factor in track or vertex reconstruction is the transverse impact parameter requirement, which is of the order of 5 mm for both experiments. Even in the most extreme cases, where the daughter particles’ tracks have a large opening angle, decays before 5 mm would yield reconstructed tracks. We therefore choose $p_{th}$ to be 5 mm. We also determine if particles decay outside of the detector, conservatively taking the dimensions of the ATLAS detector (13 m in radius, and 22 m in length on each side) to determine if a particle was undetectable (if electrically neutral).
@@
For each point of the grid, we parse the HEPMC file using PyHEPMC. For each event, based on the lifetime and Lorentz boost obtained in previous steps, we determine if each particle decayed a) promptly (with transverse decay position below $p_{th}$), b) in the detector or c) outside of the detector. For particles in case a) nothing happens: the particles are hence treated as if they were prompt. If particles decay after the prompt threshold (cases b) or c)), then the particle, and all its daughters in the decay history in the HEPMC event, are removed using the PyHEPMC functionality. The rest of the event is left intact. A final consideration is that if the decay of a neutral particle takes place within the detector, then it could be that measurements discard such events as instrumental noise. Hence, if any neutral LLPs fall in category b) and not in category c), then the entire event is filtered from the HEPMC file, though that event is still considered when calculating the total number of generated events for normalisation purposes. A new filtered HEPMC is created in this way, for each point in the parameter grid.
@@
The effect of this method is to reduce the sensitivity Contur would have reported if no filtering had taken place, since we only ever remove events and/or particles, although we note that one possible exception is if the procedure introduces excessive missing transverse energy, while little would be expected from the theory. We can in principle identify three regimes: first, when LLPs have low lifetimes, the method is effectively extrapolating the prompt sensitivity according to the proportion of reconstructible particles. On the other hand, in the long lifetime regime, LLPs are treated as missing energy, but there can be residual sensitivity to their existence, for example by observing the effect of their recoil on the transverse momentum of any prompt particles which they are produced with. In the intermediate regime, where neutral LLPs decay within the detector, we remove events entirely while conservatively assuming that such events in data could be discarded and hence not be accounted for in the existing measurements. This implies a “dip” in sensitivity from this method in the intermediate regime. Thankfully, this is exactly the regime which prompt searches are designed to probe: hence illustrating the nice complementarity of the modified Contur method we propose with the direct search programme.
@@ 4 Constraints on feebly interacting DM
To determine any additional constraints that could be extracted from existing measurements using Contur, we simulated the process encoded in the publicly available UFO file for the model FIMP_UFO using MadGraph5_aMC@NLO v3.4.2 Alwall:2014hca and Pythia v8 Sjostrand:2007gs for hadronisation. For each point, 10 000 events were generated. The grid was composed of LLP masses in the range of 100 to 1000 GeV in steps of 45 GeV. The lifetime of the $F$ was assumed to be independent of its mass, and 20 values of $c\tau$ were scanned on a logarithmic scale between 0.01 mm and 10 m. The mass of the DM candidate was set to 12 keV, although the analysis presented here is not strongly sensitive to this choice. Since the LLP $F$ is electrically charged, it would in principle form a reconstructible track before its decay. After the decay, a SM charged lepton may continue the track, but since the unseen $s$ is emitted, one would expect the track from the $F$$\ell$ to be “kinked”: in other words for the hits not to be aligned in the expected pattern for a charged particle trajectory, but instead be the discontinuous union of two curved tracks. Such signatures would be missed by ATLAS and CMS, as standard tracking algorithms look for a single particle trajectory. If the decay takes place before the prompt threshold discussed in Section 3, the lepton track would still be reconstructed as normal. We then follow the procedure set out previously to determine the exclusions from Contur. In this example, since the LLP is electrically charged, even if the decay happens outside of the bounds of the detector, the particle would not be reconstructed as missing transverse energy.
 $\gamma$  $\ell_{1}\ell_{2}$+$E_{T}^{\rm miss}$+jet ATLAS:2021jgw  $\ell_{1}\ell_{2}$+$E_{T}^{\rm miss}$ ATLAS:2019rob  $\ell$+$E_{T}^{\rm miss}$+jet
@@
In Figure 3, we compare the results obtained from Contur to limits derived from direct searches in Ref. Brooijmans:2020yij . In this case, the direct searches provide more stringent exclusions at high LLP masses, which is perhaps not surprising given that they were originally searches for high-mass supersymmetric particles, and would be more optimal in such a regime. However, all these searches have a cut-off around 150 GeV below which they do not set exclusions. This region is filled in by the Contur exclusion lines.

### v3

@@ 3 Adapting Contur to LLP topologies
We propose the following approach to treat LLPs in the Rivet and Contur machinery. The basic idea is to evaluate, for a given set of model parameters, the lab-frame decay lengths for all particles with small widths (and therefore long lifetimes) as determined by the event generator. We can then sort the particles into four categories:
  • a) those which decayed early enough to be treated as prompt;
  • b) neutral or charged particles which decayed within the detector volume and hence would be unlikely to be picked up by the standard reconstruction algorithms used for measurements if not actively discarded;
  • c) (heavy) charged particles which decayed outside the detector, but whose interactions in the detector would lead to problematic reconstruction (such as anomalous ionization) which would potentially be vetoed in most measurements;
  • d) neutral particles which decayed outside the bounds of the detector and which would be reconstructed as missing transverse energy.
The individual particles which were not reconstructible can be manually removed from the simulated events (for instance with PyHEPMC BUCKLEY2021107310 ). Events which contained decays within the detector volume or which would have anomalous interactions with the detector, causing signatures similar to instrumental noise, can be entirely removed. Once the relevant particles and/or events have been removed from the simulations, the rest of the Contur workflow can then run normally. We have implemented this workflow, and extracted new constraints on LLP models as discussed in Sections 456 and 7. This approach is conceptually related to reweighting-based approaches previously described in SModelS Heisig:2018kfq and in a standalone study of dark radiation using LLP searches Bernreuther:2022bdw , however, it goes further by making event-by-event decisions on which particles to reconstruct, and it is the first time that any such approach has been used with detector-corrected LHC measurements. We will now review the step-by-step implementation of our method, including technical details.
@@
We determine a transverse “prompt threshold” $p_{th}$ before which particles can be considered indistinguishable from prompt. Indeed, if the decay of the LLP takes place before $p_{th}$, then tracks and deposits from the decay products should be reconstructible using the standard ATLAS and CMS algorithms. To choose the value of the threshold, we are guided by Refs. ATLAS:2023nze and CMS:2014pgm . The ultimate limiting factor in track or vertex reconstruction is the transverse impact parameter requirement, which is of the order of 5 mm for both experiments. Even in the most extreme cases, where the daughter particles’ tracks have a large opening angle, decays before 5 mm would yield reconstructed tracks. We therefore choose $p_{th}$ to be 5 mm. We also determine if particles decay outside of the detector, conservatively taking the dimensions of the ATLAS detector (13 m in radius, and 22 m in length on each side).
@@
For each point of the grid, we parse the HEPMC file using PyHEPMC. For each event, based on the lifetime and Lorentz boost obtained in previous steps, we determine if each particle decayed a) promptly (with transverse decay position below $p_{th}$), b) in the detector , c) outside of the detector and charged or d) outside the detector and neutral. For particles in case a) nothing happens: the particles are hence treated as if they were prompt. If any LLP (charged or neutral) decays after the prompt threshold but within the detector, or if a charged particle decays outside the detector (cases b) and c)), the whole event is removed. Finally, in case d), the particle and all its daughters in the decay history in the HEPMC event are removed using the PyHEPMC functionality, with the rest of the event left intact. Events which are removed are still considered when calculating the total number of generated events for normalisation purposes. A new filtered HEPMC is created in this way, for each point in the parameter grid.
@@
The effect of this method is to reduce the sensitivity Contur would have reported if no filtering had taken place, since we only ever remove events and/or particles, although we note that one possible exception is if the procedure introduces excessive missing transverse energy, while little would be expected from the theory. We can in principle identify three regimes: first, when LLPs have low lifetimes, the method is effectively extrapolating the prompt sensitivity according to the proportion of reconstructible particles. On the other hand, in the long lifetime regime, neutral LLPs are treated as missing energy, but there can be residual sensitivity to their existence, for example by observing the effect of their recoil on the transverse momentum of any prompt particles which they are produced with. In the intermediate regime, where LLPs decay within the detector, we remove events entirely while conservatively assuming that such events in data could be discarded and hence not be accounted for in the existing measurements. This implies a “dip” in sensitivity from this method in the intermediate regime for neutral LLPs. Thankfully, this is exactly the regime which direct LLP searches are designed to probe: hence illustrating the nice complementarity of the modified Contur method we propose with the direct search programme.
@@
The method outlined in this section has a potential downside. Indeed, in regimes where many events are removed, the results can be subject to large statistical fluctuations. This effect is mitigated in the results which follow by the fact that the statistical uncertainties of the signal samples are accounted for in the Contur machinery when calculating limits. Further, the most stringent (and relevant) limits are obtained in regions where the fraction of kept events is high. However, we note that it is in principle possible to calculate the probability of an event being kept analytically rather than relying on sampling an exponential. We leave this development for future work.
@@ 4 Constraints on feebly interacting DM
To determine any additional constraints that could be extracted from existing measurements using Contur, we simulated the process encoded in the publicly available UFO file for the model FIMP_UFO using MadGraph5_aMC@NLO v3.4.2 Alwall:2014hca and Pythia v8 Sjostrand:2007gs for hadronisation. For each point, 10 000 events were generated. The grid was composed of LLP masses in the range of 100 to 1000 GeV in steps of 45 GeV. The lifetime of the $F$ was assumed to be independent of its mass, and 20 values of $c\tau$ were scanned on a logarithmic scale between 0.01 mm and 10 m. The mass of the DM candidate was set to 12 keV, although the analysis presented here is not strongly sensitive to this choice. Since the LLP $F$ is electrically charged, it would in principle form a reconstructible track before its decay, although this could have anomalous ionization and hence might not be reconstructed properly. After the decay, a SM charged lepton may continue the track, but since the unseen $s$ is emitted, one would expect the track from the $F$$\ell$ to be “kinked”: in other words for the hits not to be aligned in the expected pattern for a charged particle trajectory, but instead be the discontinuous union of two curved tracks. Such signatures would be missed by ATLAS and CMS, as standard tracking algorithms look for a single particle trajectory. Both these consideration motivate the need to veto events where the decay takes place after the prompt threshold. If the decay takes place before the prompt threshold discussed in Section 3, the lepton track would still be reconstructed as normal. We then follow the procedure set out previously to determine the exclusions from Contur.
  No data  $\ell_{1}\ell_{2}$+$E_{T}^{\rm miss}$+jet ATLAS:2021jgw  $\ell_{1}\ell_{2}$+$E_{T}^{\rm miss}$ ATLAS:2019rob  $\ell$+$E_{T}^{\rm miss}$+jet
@@
In Figure 3, we compare the results obtained from Contur to limits derived from direct searches in Ref. Brooijmans:2020yij . In this case, the direct searches provide more stringent exclusions at high LLP masses, which is perhaps not surprising given that they were originally searches for high-mass supersymmetric particles, and would be more optimal in such a regime. However, these searches have a cut-off around 150 GeV below which they do not set exclusions, although we note that this region is filled in by 8 TeV direct searches for prompt particles ATLAS:2018ojr ; ATLAS:2014zve . That region is also covered by the Contur exclusion lines.